OpenStudy (anonymous):

Evaluate and, if possible, simplify 18/-9 -x for x=-9

OpenStudy (anonymous):

18/-9=-2-(-9)=7

OpenStudy (lgbasallote):

if it's \(\Large \frac{-18}{-9 - x}\) then it's undefined lol

OpenStudy (anonymous):

that second reply is correct if it -9-(-9) is the whole denominator! sorry about that!